OpenStudy (dancergirl45):

When the function f(x) = 6(9)x is changed to f(x) = 6(9)x + 1, what is the effect? There is no change to the graph because the exponential portion of the function remains the same. All input values are moved 1 space to the right. The x-intercept is 1 space higher. The y-intercept is 1 space higher.

OpenStudy (dancergirl45):

@welshfella

OpenStudy (dancergirl45):

I think c

OpenStudy (tkhunny):

1) Don't post problems with no work shown. 2) 6(9)x means 6*9*x. Did you mean 6*9^x = \(6\cdot 9^{x}\)? 3) Where does the "+1" go? \(6\cdot 9^{x+1}\;or\;6\cdot 9^{x}+1\)

OpenStudy (dancergirl45):

@pooja195

OpenStudy (danjs):

graph y = A*B^(x - C) + D https://www.desmos.com/calculator hit add slider button for all, play with each one A B C D , see how it changes the graph

OpenStudy (dancergirl45):

where is the add slider button

OpenStudy (danjs):

make B a little larger than 1

OpenStudy (danjs):

after you type in the equation it will show the buttons below it

OpenStudy (dancergirl45):

so it is C

OpenStudy (dancergirl45):

or no D

OpenStudy (dancergirl45):

D right

OpenStudy (danjs):

yeah the whole graph shifts in ydirection y = A*B^(x - C) + D D has no part of what the x term does , it is just an added on constant , every Y value will be changed by D

OpenStudy (dancergirl45):

So D

OpenStudy (danjs):

yes.

OpenStudy (dancergirl45):

Thank you so much!! you are so helpful, can you help me with another

OpenStudy (danjs):

C - xdirection shifts, any x you had is now C different A - scales the exponent, justmultiplies the value by an amount , you can see the effects on the graph, stretched upwards

OpenStudy (danjs):

sure, one more